2. The light reflected from a soap bubble (n-1.35) appears green( What is the minimum thickness (in nm)? 540 nm) at its center. (a) 55 (b) 100 (c) 115 (d) 135 (e) 205
Answer #1

2 \mu. t = n \lambda

2*1.35*t = 1 * 540*10-9

t = 200 nm

minimum thickness = 200/2